OpenStudy (yusmin):

y=-x+2 (find the Y intercept ) type in order pair?? please help

OpenStudy (yusmin):

oops its y=-4x+2

OpenStudy (oswaldmurphy):

I am not exactly sure what you are asking but, y = -1x + 2. y = mx+b. b= y-int and m=slope.

OpenStudy (oswaldmurphy):

Oh ok, y = -4x + 2. You are trying to find the ordered pair for the y-intercept correct?

OpenStudy (yusmin):

yes

OpenStudy (oswaldmurphy):

Well, rise over run. I think what you would do is just simply get (2,0). Since you start off going up 2 since it is a positive.

OpenStudy (yusmin):

and the slope would be -4 right?

jimthompson5910 (jim_thompson5910):

@OswaldMurphy you have the coordinates mixed up. It should be (0,2) x comes before y

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):